# Testing Seesaw and Leptogenesis via Gravitational Waves: Majorana versus Dirac

### v1

@@ III Dirac Neutrinos and Lepton Number Violating Dirac Leptogenesis
In Fig. 2 , we show the vertex and self-energy loop diagrams that contribute to the non-zero CP-violation. The decay of the lightest $\zeta_{i}$ leads to a CP-asymmetry Heeck:2013vha ,
@@ IV Primordial GW spectrum from Local Cosmic Strings
Cosmic Strings (CS) may originate as fundamental objects such as in String Theory Copeland:2009ga ; Polchinski:2004ia as well as topological defects in field theory during a symmetry-breaking phase transition. We will focus on cosmic strings that were formed in the early Universe due to the spontaneous breaking of a local $U(1)$ symmetry, this is sometimes called field theory strings. The simplest model where cosmic strings cmay arise is a $U(1)$ local gauge or global theory in the context of a complex scalar field denoted by $\varphi$. The Lagrangian density for such a theory can be written as,
@@
$V(\varphi)=\frac{1}{2}\lambda(\varphi^{*}\varphi-\frac{1}{2}\eta^{2})^{2}$ (17)
where $\lambda$ is the self-quartic coupling and $\langle\varphi\rangle=\eta$ is the vacuum expectation value (VEV) of the scalar field.
@@
Cosmic strings are essentially field configurations at the top of this Mexican hat potential. The CS have a negligible thickness (with widths given by approximately $1/(\sqrt{\lambda}\eta$)) and hence their dynamics in an expanding Universe are well described by something known as the Nambu-Goto action, in its zero-width approximation. The mass per unit length $\mu$ of local CS is given by Gouttenoire:2019kij ,
@@
where $\eta$ is the vacuum expectation value of the scalar field constituting the CS and n is the (integer-valued) winding number of the cosmic string formed. Only $n=1$ is considered to be a stable configuration Bettencourt:1996qe ; Ghoshal:2023sfa . The local CS network starts forming at a temperature,
@@
Straight infinitely long strings are stable against decay due to their topological nature, therefore they contribute negligibly towards the GW spectra compared to cosmic string loops. Local CS loops, formed due to intercommutation events of CS, are free of any topological charge, so they contribute significantly towards GW radiation emission Vilenkin:2000jqa . The GW radiation power for CS loops is given by,
@@
where $\Gamma$ is the total GW emission efficiency of loops and is determined typically from the Nambu-Goto simulation. This has been estimated to be $\Gamma\simeq 70$ Blanco-Pillado:2013qja .
The GW spectrum from Local CS Loops observed today is given follows Gouttenoire:2019kij :
@@
and $t_{\mathrm{osc}}=\operatorname{Max}\left[t_{fric},t_{F}\right]$, where $t_{F}$ is the time of CS network formation, defined as $\sqrt{\rho_{tot}\left(t_{F}\right)}\equiv\mu$ where $\rho_{tot}$ is the total energy budget of the Universe. In the presence of friction, the string motion is damped at high temperature until the time $t_{fric}$. $dn_{\mathrm{loop}}/dt_{i}$ expresses the rate of formation of loops with a distribution size $\mathcal{P}_{\mathrm{loop}}(\alpha)$. These objects initially redshift as $a^{-3}$ before radiating GW with power $P_{\mathrm{GW}}^{(k)}$ after which they dilute as $a^{-4}$. Using such simplifications, one may express Eq. 22 as Gouttenoire:2019kij :
@@
where $\Omega_{r}h^{2}\simeq 4.2\times 10^{-5}$ ParticleDataGroup:2020ssz . $\tilde{T_{M}}$ is the temperature of the Universe at the time of the maximal GW emission $\tilde{t_{M}}$,
@@ V Numerical results
    Ground based interferometers: Laser Interferometer Gravitational-wave Observatory (LIGO) LIGOScientific:2016aoc ; LIGOScientific:2016sjg ; LIGOScientific:2017bnn ; LIGOScientific:2017vox ; LIGOScientific:2017ycc ; LIGOScientific:2017vwq , Advanced LIGO (a-LIGO) LIGOScientific:2014pky ; LIGOScientific:2019lzm , Einstein Telescope (ET) Punturo_2010 ; Hild:2010id , Cosmic Explorer (CE) Reitze:2019iox .
@@
    Space based interferometers: $\mu$-ARES Sesana:2019vho , Laser Interferometer Space Antenna (LISA) amaroseoane2017laser ; Baker:2019nia , Big-Bang Observer (BBO) Corbin:2005ny ; Harry_2006 , Deci-Hertz Interferometer Gravitaitonal-wave Observatory (DECIGO) Yagi:2011yu , Upgraded DECIGO (U-DECIGO) Seto:2001qf ; Kawamura_2006 ; Yagi:2011wg .
@@
    Pulsar Timing Arrays (PTA): European Pulsar Timing Array These are recast of star surveys (EPTA) Kramer:2013kea ; Lentati:2015qwp ; Babak:2015lua , Square Kilometre Array (SKA) Janssen:2014dka ; Weltman:2018zrl ; Carilli:2004nx , North American Nanohertz Observatory for Gravitational Waves (NANOGrav) McLaughlin:2013ira ; NANOGRAV:2018hou ; Aggarwal:2018mgp ; Brazier:2019mmu ; NANOGrav:2020bcs .
@@ V.3 Comparison: Dirac versus Majorana Leptogenesis
In general, the BLB-L number is violated in Majorana Leptogenesis, whereas in Dirac Leptogenesis the BLB-L is conserved Dick:1999je ; Narendra:2017uxl . In current scenario, the Dirac Leptogenesis the BLB-L number violates similar to the Majorana scenario Heeck:2013vha . The origin of lepton asymmetry is different in both scenarios. In Dirac Leptogenesis case a heavy scalar generates asymmetry in right-handed neutrinos and then transfer to the lepton asymmetry, whereas in Majorana Leptogenesis case a heavy right-handed neutrino generates the lepton asymmetry. In the following, to establish a comparison between these scenarios we compare the mass scales of the Leptogenesis.
@@ VI Discussion and Conclusion
Thermal Dirac Leptogenesis through the Dirac seesaw mechanism gives an elegant and minimal explanation for two outstanding puzzles in the standard model namely the generation of tiny neutrino masses and matter-antimatter asymmetry of the Universe involving no lepton number violation (LNV) on the whole [^10] In B-L extended Dirac leptogenesis one does have LNV.. Unfortunately, the scales of new physics, namely the seesaw scale and the Leptogenesis scale are naturally well beyond what we can directly test in terrestrial laboratory experiments. Given the fundamental nature of these puzzles, indirect cosmological tests of Dirac Leptogenesis are of great value and we propose cosmic strings as a powerful witness of the paradigm. We find that, if present, gravitational wave radiation from cosmic strings can probe all the parameter space relevant for thermal Dirac Leptogenesis. If there is U(1)B-L survives below from GUT symmetry breaking scales to these scales to protect the right handed neutrino mass, then we show that cosmic strings appear through the breaking of this B-L symmetry predicting a spectrum of stochastic gravitational waves and, with future detectors being able to probe the entire mass range relevant to the paradigm of Dirac Leptogenesis. If cosmic strings GW spectrum is discovered and falls into the energy scales relevant for the Dirac seesaw mechanism and Leptogenesis, it would provide exciting hints of dynamics in the lepton sector at high energy scales. We showed a comparative analysis between such probes of GW in the context of Majorana and Dirac Leptogenesis within the B-L extension paradigm, and based on the energy scales of seesaw and Leptogenesis. We have shown, for instance, LISA detector will be able to probe the scale of seesaw $\langle\varphi\rangle=10^{11}$ GeV or higher while ET will be able to probe $\langle\varphi\rangle=10^{12}$ GeV or higher. Along with these, Figs. 6 and 7 respectively also tell us about the scales of Dirac and Majorana seesaws, for which correct BAU can be satisfied by varying the Yukawas ($\text{Tr}[\lambda^{\dagger}\lambda]$) which control the CP-violation. It can be inferred that for the Majorana case, $\langle\varphi\rangle\geq 10^{12}$ GeV agree with observed BAU while below $\langle\varphi\rangle\leq 10^{12}$ GeV, the $Y_{B}\leq 8.66\times 10^{-11}$ given $y_{m}=1,\frac{M_{\rm N_{2}}}{M_{\rm N_{1}}}=10^{2}$. This situation is however relaxed for the Dirac seesaw case, and the observed BAU can be satisfied for a wide range of values of $\langle\varphi\rangle$. Given the correlations between the scale of seesaw and scale of Leptogenesis, which are different for Dirac Leptogenesis and Majorana Leptogenesis, ET detector will be able to probe the scale of Dirac Leptogenesis $M_{\rm\zeta_{1}}\sim 10^{9}$ GeV or higher (see Fig. 8 ), while the same for Majorana Leptogenesis $M_{N_{1}}\geq 10^{12}$ GeV within perturbative coupling limits (see Fig. 9 ). LISA on the other hand will be able to probe even upto lower scales of Dirac Leptogenesis, $M_{\rm\zeta_{1}}\sim 5\times 10^{8}$ GeV, and that of Majorana Leptogenesis $M_{N_{1}}\gtrsim 8\times 10^{11}$ GeV.
@@ Appendix A Appendix: Signal-to-noise ratio (SNR)
Gravitational Wave Interferometers actually measure displacements of detector arms in terms of a what is known in terms of dimensionless strain-noise $h_{\text{GW}}(f)$ which is related to the GW amplitude. This can be straight-forwardly converted into the corresponding energy density Garcia-Bellido:2021zgu
@@
where $H_{0}=h\times 100\;\text{(km/s)}/\text{Mpc}$ is the Hubble expansion rate today. The signal-to-noise ratio (SNR) for a projected experimental sensitivity $\Omega_{\text{exp}}(f)h^{2}$ is then estimated in order to assess the detection probability of the primordial GW background originating from the global cosmic string background following the prescription Thrane:2013oya ; Caprini:2015zlo

### v2

@@ III Dirac Neutrinos and Lepton Number Violating Dirac Leptogenesis
In Fig. 2 , we show the vertex and self-energy loop diagrams that contribute to the non-zero CP-violation. The decay of the lightest $\zeta_{i}$ leads to a CP-asymmetry Heeck (2013),
@@ IV Primordial GW spectrum from Local Cosmic Strings
Cosmic Strings (CS) may originate as fundamental objects such as in String Theory Copeland and Kibble (2010); Polchinski (2004) as well as topological defects in field theory during a symmetry-breaking phase transition. We will focus on cosmic strings that were formed in the early Universe due to the spontaneous breaking of a local $U(1)$ symmetry, this is sometimes called field theory strings. The simplest model where cosmic strings cmay arise is a $U(1)$ local gauge or global theory in the context of a complex scalar field denoted by $\varphi$. The Lagrangian density for such a theory can be written as,
@@
$V(\varphi)=\frac{1}{2}\lambda^{\prime}(\varphi^{*}\varphi-\frac{1}{2}\eta^{2})^{2}$ (17)
where $\lambda^{\prime}$ is the self-quartic coupling and $\langle\varphi\rangle=\eta$ is the vacuum expectation value (VEV) of the scalar field.
@@
Cosmic strings are essentially field configurations at the top of this Mexican hat potential. The CS have a negligible thickness (with widths given by approximately $1/(\sqrt{\lambda}\eta$)) and hence their dynamics in an expanding Universe are well described by something known as the Nambu-Goto action, in its zero-width approximation. The mass per unit length $\mu$ of local CS is given by Gouttenoire et al. (2020),
@@
where $\eta$ is the vacuum expectation value of the scalar field constituting the CS and n is the (integer-valued) winding number of the cosmic string formed. Only $n=1$ is considered to be a stable configuration Bettencourt et al. (1997); Ghoshal et al. (2023a). The local CS network starts forming at a temperature,
@@
Straight infinitely long strings are stable against decay due to their topological nature, therefore they contribute negligibly towards the GW spectra compared to cosmic string loops. Local CS loops, formed due to intercommutation events of CS, are free of any topological charge, so they contribute significantly towards GW radiation emission Vilenkin and Shellard (2000). The GW radiation power for CS loops is given by,
@@
where $\Gamma$ is the total GW emission efficiency of loops and is determined typically from the Nambu-Goto simulation. This has been estimated to be $\Gamma\simeq 70$ Blanco-Pillado et al. (2014).
The GW spectrum from Local CS Loops observed today is given follows Gouttenoire et al. (2020):
@@
and $t_{\mathrm{osc}}=\operatorname{Max}\left[t_{fric},t_{F}\right]$, where $t_{F}$ is the time of CS network formation, defined as $\sqrt{\rho_{tot}\left(t_{F}\right)}\equiv\mu$ where $\rho_{tot}$ is the total energy budget of the Universe. In the presence of friction, the string motion is damped at high temperature until the time $t_{fric}$. $dn_{\mathrm{loop}}/dt_{i}$ expresses the rate of formation of loops with a distribution size $\mathcal{P}_{\mathrm{loop}}(\alpha)$. These objects initially redshift as $a^{-3}$ before radiating GW with power $P_{\mathrm{GW}}^{(k)}$ after which they dilute as $a^{-4}$. Using such simplifications, one may express Eq. 22 as Gouttenoire et al. (2020):
@@
where $\Omega_{r}h^{2}\simeq 4.2\times 10^{-5}$ Zyla and others (2020a). $\tilde{T_{M}}$ is the temperature of the Universe at the time of the maximal GW emission $\tilde{t_{M}}$,
@@ V Numerical results
    Ground based interferometers: Laser Interferometer Gravitational-wave Observatory (LIGO) Abbott and others (2016b, a, 2017d, 2017a, 2017b, 2017c), Advanced LIGO (a-LIGO) Aasi and others (2015); Abbott and others (2021), Einstein Telescope (ET) M. Punturo (2010); Hild and others (2011), Cosmic Explorer (CE) Reitze and others (2019).
@@
    Space based interferometers: $\mu$-ARES Sesana and others (2021), Laser Interferometer Space Antenna (LISA) Amaro-Seoane et al. (2017); Baker and others (2019), Big-Bang Observer (BBO) Corbin and Cornish (2006); Harry et al. (2006), Deci-Hertz Interferometer Gravitaitonal-wave Observatory (DECIGO) Yagi et al. (2011), Upgraded DECIGO (U-DECIGO) Seto et al. (2001); Kawamura et al. (2006); Yagi and Seto (2011).
@@
    Pulsar Timing Arrays (PTA): European Pulsar Timing Array These are recast of star surveys (EPTA) Kramer and Champion (2013); Lentati and others (2015b); Babak and others (2016a), Square Kilometre Array (SKA) Janssen and others (2015); Weltman and others (2020); Carilli and Rawlings (2004), North American Nanohertz Observatory for Gravitational Waves (NANOGrav) McLaughlin (2013); Arzoumanian and others (2018); Aggarwal and others (2019); Brazier and others (2019); Arzoumanian and others (2020).
@@ V.3 Comparison: Dirac versus Majorana Leptogenesis
In general, the BLB-L number is violated in Majorana Leptogenesis, whereas in Dirac Leptogenesis the BLB-L is conserved Dick et al. (2000); Narendra et al. (2018b). In current scenario, the Dirac Leptogenesis the BLB-L number violates similar to the Majorana scenario Heeck (2013). The origin of lepton asymmetry is different in both scenarios. In Dirac Leptogenesis case a heavy scalar generates asymmetry in right-handed neutrinos and then transfer to the lepton asymmetry, whereas in Majorana Leptogenesis case a heavy right-handed neutrino generates the lepton asymmetry. In the following, to establish a comparison between these scenarios we compare the mass scales of the Leptogenesis.
@@ VI Discussion and Conclusion
Thermal Dirac Leptogenesis through the Dirac seesaw mechanism gives an elegant and minimal explanation for two outstanding puzzles in the standard model namely the generation of tiny neutrino masses and matter-antimatter asymmetry of the Universe involving no lepton number violation (LNV) on the whole  [^14] In B-L extended Dirac leptogenesis one may have LNV.. Unfortunately, the scales of new physics, namely the seesaw scale and the Leptogenesis scale are naturally well beyond what we can directly test in terrestrial laboratory experiments. Given the fundamental nature of these puzzles, indirect cosmological tests of Dirac Leptogenesis are of great value and we propose cosmic strings as a powerful witness of the paradigm. We find that, if present, gravitational wave radiation from cosmic strings will carry the imprints of thermal Dirac leptogenesis. If there is U(1)B-L survives below from GUT symmetry breaking scales to these scales to protect the right handed neutrino mass, then we show that cosmic strings appear through the breaking of this B-L symmetry predicting a spectrum of stochastic gravitational waves and, with future detectors being able to probe the entire mass range relevant to the paradigm of Dirac Leptogenesis. If cosmic strings GW spectrum is discovered and falls into the energy scales relevant for the Dirac seesaw mechanism and Leptogenesis, it would provide exciting hints of dynamics in the lepton sector at high energy scales. We showed a comparative analysis between such probes of GW in the context of Majorana and Dirac Leptogenesis within the B-L extension paradigm, and based on the energy scales of seesaw and Leptogenesis. We have shown, for instance, LISA detector will be able to probe the scale of seesaw $\langle\varphi\rangle=10^{11}$ GeV or higher while ET will be able to probe $\langle\varphi\rangle=10^{12}$ GeV or higher. Along with these, Figs. 6 and 7 respectively also tell us about the scales of Dirac and Majorana seesaws, for which correct BAU can be satisfied by varying the Yukawas ($\text{Tr}[\lambda^{\dagger}\lambda]$) which control the CP-violation. It can be inferred that for the Majorana case, $\langle\varphi\rangle\geq 10^{12}$ GeV agree with observed BAU while below $\langle\varphi\rangle\leq 10^{12}$ GeV, the $Y_{B}\leq 8.66\times 10^{-11}$ given $y_{m}=1,\frac{M_{\rm N_{2}}}{M_{\rm N_{1}}}=10^{2}$. This situation is however relaxed for the Dirac seesaw case, and the observed BAU can be satisfied for a wide range of values of $\langle\varphi\rangle$. Given the correlations between the scale of seesaw and scale of Leptogenesis, which are different for Dirac Leptogenesis and Majorana Leptogenesis, ET detector will be able to probe the scale of Dirac Leptogenesis $M_{\rm\zeta_{1}}\sim 10^{9}$ GeV or higher (see Fig. 8 ), while the same for Majorana Leptogenesis $M_{N_{1}}\geq 10^{12}$ GeV within perturbative coupling limits (see Fig. 9 ). LISA on the other hand will be able to probe even upto lower scales of Dirac Leptogenesis, $M_{\rm\zeta_{1}}\sim 5\times 10^{8}$ GeV, and that of Majorana Leptogenesis $M_{N_{1}}\gtrsim 8\times 10^{11}$ GeV.
@@ Appendix A Appendix: Signal-to-noise ratio (SNR)
Gravitational Wave Interferometers actually measure displacements of detector arms in terms of a what is known in terms of dimensionless strain-noise $h_{\text{GW}}(f)$ which is related to the GW amplitude. This can be straight-forwardly converted into the corresponding energy density Garcia-Bellido et al. (2021)
@@
where $H_{0}=h\times 100\;\text{(km/s)}/\text{Mpc}$ is the Hubble expansion rate today. The signal-to-noise ratio (SNR) for a projected experimental sensitivity $\Omega_{\text{exp}}(f)h^{2}$ is then estimated in order to assess the detection probability of the primordial GW background originating from the global cosmic string background following the prescription Thrane and Romano (2013); Caprini and others (2016)